OpenStudy (anonymous):

Please help me! I'll fan and give a medal! The function f(x)=16(5)^x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@jhonyy9

OpenStudy (phi):

this one is a bit tricky.

OpenStudy (anonymous):

How so?

OpenStudy (solomonzelman):

2 times a year, then make an exponent 2x. In other words multiply the exponent times 2.

OpenStudy (anonymous):

I thought that may be it- is there any reason to change the number in the parentheses, or the coeffienct?

OpenStudy (solomonzelman):

I am assuming that every growth, is different, and that 2 growths in one year aren;t same.

OpenStudy (anonymous):

I think so- I'm not entirely sure. The problem is kind of vague regarding that...

OpenStudy (solomonzelman):

Or if the 2 growths in the year are same, then make it 10, instead of 5.

OpenStudy (anonymous):

Well, that isn't in any of the answer choices, so let's assume that they vary

OpenStudy (solomonzelman):

So, if the 2 growths in a year vary, you will get \(\normalsize\color{black}{ f(x)=16(5)^{2x}}\), and if they don't then you get \(\normalsize\color{black}{ f(x)=16(10)^{x}}\).\(\LARGE\color{white}{ \rm │ }\)

OpenStudy (anonymous):

Ok- how would you go about finding the new growth rate?

OpenStudy (anonymous):

just solve for x?

OpenStudy (phi):

Here is how to do it: \[ 16(5)^x = 16(1+ 4)^x \] if we double to two per year, we use 2x as the exponent. However, the growth rate (the 4) must change (call it y): \[ 16(1+ y)^{2x} \] we want that to give the same number as the original equation. in other words, after 1 year: 16(5)^1 gives 16*5 we want \[ 16(1+ y)^{2} = 16 \cdot 5\\ (1+ y)^{2} = 5 \\ 1+y= \sqrt{5} \\ y= \sqrt{5}-1\]

OpenStudy (phi):

the answer will be \[ 16(1+\sqrt{5}-1 )^{2x} \\ 16(2.24)^{2x} \]

OpenStudy (anonymous):

That makes sense- thank you @phi and @SolomonZelman

OpenStudy (solomonzelman):

with 2 growths the only thing changes that there are 2 growths, but not the overall growth per year ?

OpenStudy (anonymous):

I'm not sure what you mean

OpenStudy (solomonzelman):

I mean that he is just basically doing this, \(\normalsize\color{black}{ f(x)=16(5)^x}\) is equivalent to, \(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ f(x)=16(\sqrt{5}^2)^x}\)\(\LARGE\color{white}{ \rm │ }\) \(\normalsize\color{black}{ f(x)=16(\sqrt{5})^{2x}}\)\(\LARGE\color{white}{ \rm │ }\)

OpenStudy (solomonzelman):

But the growth per year doesn't change from this

OpenStudy (solomonzelman):

when it should

OpenStudy (solomonzelman):

I am not trying to argue, I am just trying to understand this...

OpenStudy (anonymous):

Ok... I don't mean to be dense, but isn't that essentially what you did? xcept he added the radical sign in the parentheses?

OpenStudy (phi):

Here is a graph of both functions. As you see, they are identical curves. However, they have different growth rates.

OpenStudy (solomonzelman):

I see, so it is just that you do not change the yearly growth... then yes

OpenStudy (anonymous):

So, the growth rate is exactly the same, just written differently?

OpenStudy (anonymous):

oi

OpenStudy (solomonzelman):

I thought you change the growth, adding another, additional, growth `×5` so that there are two `×5` growths..... (misinterpretation )

OpenStudy (solomonzelman):

the growth rate is different as phi showed.

OpenStudy (solomonzelman):

I am just saying what I thought the problem was

OpenStudy (anonymous):

Ok! Thank @SolomonZelman

OpenStudy (anonymous):

*thanks

OpenStudy (anonymous):

Would you mind helping me with a problem an arithmetic sequence?

OpenStudy (solomonzelman):

Anytime... even though I didn't actually help.

OpenStudy (solomonzelman):

Yes, if I know how to do it....

OpenStudy (anonymous):

I thought you did :)

OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 6 and a second term of 2?

OpenStudy (solomonzelman):

domain of n, or the terms ?

OpenStudy (phi):

thinking about it, I made it too complicated. To solve do this \[ 16(5)^x = 16(5)^{\frac{1}{2} \cdot 2 \cdot x} \\ = 16(5^\frac{1}{2})^{2x} \\ = 16 (\sqrt{5})^{2x} \]

OpenStudy (solomonzelman):

\(\large\color{black}{ a_1}\) is 6, and \(\large\color{black}{ a_2}\) is 2,. So d= ?

OpenStudy (solomonzelman):

can you find the difference ?

OpenStudy (anonymous):

I think so- the answer choices show the explicit equation, then an inequality of n and 1 or 0 (they vary)

OpenStudy (anonymous):

4?

OpenStudy (solomonzelman):

Close... can you post the choices?

OpenStudy (solomonzelman):

You subtract 4, because to get from 6 to 2, you subtract 4. Right? so d=-4.

OpenStudy (anonymous):

Sure- give me a moment- I need to type them out and I'm not exactly fast

OpenStudy (solomonzelman):

Okay :)

OpenStudy (anonymous):

a sub(n) = 6-2(n-1); all integers where n is greater than or equal to 1 a sub(n) = 6-2(n-1); all integers where n is greater than or equal to 0 a sub(n) = 6-4(n-1); all integers where n is greater than or equal to 0 a sub(n) = 6-4(n-1); all integers where n is greater than or equal to 1

OpenStudy (anonymous):

I'm ready for something big here @SolomonZelman ;)

OpenStudy (solomonzelman):

\(\large \color{black}{ a_{n}=6-2(n-1).~~~~\rm ~ all~~integers,~~ n≥1}\) \(\LARGE\color{white}{ \rm │ }\) \(\large \color{black}{ a_{n}=6-2(n-1).~~~~\rm ~ all~~integers,~~ n≥0}\) \(\LARGE\color{white}{ \rm │ }\) \(\large \color{black}{ a_{n}=6-4(n-1).~~~~\rm ~ all~~integers,~~ n≥0}\) \(\LARGE\color{white}{ \rm │ }\) \(\large \color{black}{ a_{n}=6-4(n-1).~~~~\rm ~ all~~integers,~~ n≥1}\) \(\LARGE\color{white}{ \rm │ }\)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

exactly

OpenStudy (solomonzelman):

The number in front of the `(n-1)` is the difference. Based on that, exclude A and B and the domain for n can be 1 or greater, since a term can be \(\normalsize\color{black}{ 1}\)st, \(\normalsize\color{black}{ 2}\)nd, \(\normalsize\color{black}{3}\)rd... and on. But, not \(\normalsize\color{black}{ 0}\)th. Exclude C

OpenStudy (anonymous):

So D is my correct answer?

OpenStudy (anonymous):

For all the time we spent typing the answers, that took about 10 seconds XD

OpenStudy (anonymous):

oi... I have so many more problems. This is a sad day.

OpenStudy (solomonzelman):

But at least we got this prob :) (sorry for a late reply)

OpenStudy (anonymous):

It's fine :) Thank you so much for the help! This really is a very generous thing to do

OpenStudy (solomonzelman):

yw